OpenStudy (anonymous):

Please help! Simplify the following expression: x - 4y + 5z - 3x - y + 2z = what and how? 1) -2x - 5y - 7z 2) -2x - 5y + 7z 3) -2x -5y + 3z

OpenStudy (anonymous):

I could not open my last post

OpenStudy (anonymous):

Sorry @***[ISURU]*** for trouble

OpenStudy (anonymous):

so simple x - 4y + 5z - 3x - y + 2z = x - 3x - 4y - y + 5z + 2z = - 2x - 5y + 7z